Two vectors have magnitudes 3 unit and 4 unit respectively. What should be the angle between them if the magnitude of the resultant is (a) 1 unit, (b) 5 unit and (c) 7 unit.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. |a ⃗ | = 3m |b ⃗ | = 4 (a) if R = 1 unit ⇒ √(3^2+4^2+2.3.4.cos⁡θ ) = 1 Θ = 180° (b) √(3^2+4^(2 )+22.3.4.cos⁡θ ) = 5 Θ = 90° (c) √(3^2+4^2+2.3.4.cos⁡θ ) = 7 θ = 0° Angle between them is 0°.